# Objects May Be Closer Than They Appear: Significant Host Galaxy Dispersion Measures of Fast Radio Bursts in Zoom-in Simulations

### v1

@@ Abstract
We investigate the contribution of host galaxies to the overall Dispersion Measures (DMs) for Fast Radio Bursts (FRBs) using the Feedback in Realistic Environments (FIRE-2) cosmological zoom-in simulation suite. We calculate DMs from every star particle in the simulated L* galaxies by ray-tracing through their multi-phase interstellar medium (ISM), summing the line-of-sight free thermal electron column for all gas elements within 20 kpc of the galactic mid-plane. At , we find average (median) host-galaxy DMs of 74 (43) and 210 (94) pc cm-3 for older (10 Myr) and younger (10 Myr) stellar populations, respectively. Inclination raises the median DM measured for older populations (10 Myr) in the simulations by a factor of 2, but generally does not affect the younger stars deeply embedded in HII regions except in extreme edge-on cases (inclination ). In kinematically disturbed snapshots ( in FIRE), the average (median) host-galaxy DMs are higher: 80 (107) and 266 (795) pc cm-3 for older (10 Myr) and younger (10 Myr) stellar populations, respectively. FIRE galaxies tend to have higher DM values than cosmological simulations such as IllustrisTNG. As a result, FRB host galaxies may be closer (lower redshift) than previously inferred. Furthermore, constraining host-galaxy DM distributions may help significantly constrain FRB progenitor models.
@@ 1 Introduction
First discovered by Lorimer et al. (2007), Fast Radio Bursts (FRBs) are a mysterious class of millisecond duration radio transients that are likely extragalactic in origin. To date, hundreds of FRBs have been identified and a few tens have localized redshifts through the combined efforts of multiple teams using ASKAP, Parks, MOST, CHIME, DSA-110, and FAST telescopes (e.g., Thornton et al., 2013; Caleb et al., 2017; Bhandari et al., 2018; Shannon et al., 2018; CHIME/FRB Collaboration et al., 2019a; Heintz et al., 2020; Niu et al., 2021; The CHIME/FRB Collaboration et al., 2021; Law et al., 2023). Additionally, some FRBs are known to be repeaters, providing some clues about their origins (e.g., Spitler et al., 2016; CHIME/FRB Collaboration et al., 2019a, b; Fonseca et al., 2020; Bhardwaj et al., 2021). Despite this observational progress, the origins of FRBs remain mysterious (Cordes & Chatterjee, 2019; Petroff et al., 2019; Platts et al., 2019; Zhang, 2020; Petroff et al., 2021)111So much so that popular culture has picked up on the scientific debate, e.g., see: https://xkcd.com/2886/. One promising model for FRBs is that they are associated with magnetars; an idea which is strengthened by the detection of an FRB associated with the Milky Way (MW) magnetar SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020; Lu et al., 2020; Margalit et al., 2020). However, given its low energy compared to other FRBs, other source objects are certainly still viable (Cordes & Chatterjee, 2019; Petroff et al., 2019; Platts et al., 2019; Zhang, 2020). Ultimately, to determine the origins and nature of FRBs, their host galaxy and local environmental conditions must be understood, which requires more precise localization and understanding of the signal characteristics (Li et al., 2023; Mo et al., 2023).
@@
†Star formation rates averaged over a 40 Myr period, ignoring stellar evolutionary mass loss. |
@@
However, there is some evidence that FRBs have significantly larger host-galaxy DMs than suggested by cosmological simulations. For example, repeating FRB 20190520B has an estimated host-galaxy pc cm-3 (after accounting for foreground galaxy clusters) and a redshift of (Niu et al., 2022; Ocker et al., 2022; Anna-Thomas et al., 2022; Lee et al., 2023), which provides a cautionary example of interpreting high-DM FRBs as having high redshifts.
@@ 2.2 Observed FRB Dispersion Measures
To compare our calculated host-galaxy DMs against the observed distribution of FRB DMs, we searched the Transient Name Server (TNS, https://www.wis-tns.org/) for reported FRB events. As of April 17, 2024, 43 FRBs with measured DMs (), localized redshifts, and estimates for Milky Way foreground DM () were found on the server, which we took as a representative population. These originated from TNS FRB Report (FR) 2020 ID #2470, TNS FRB Report (FR) 2021 ID #3302, and TNS FRB Report (FR) 2023 ID #419. We have not recalibrated or scaled the values taken from TNS in any way for our analysis: they are taken at face value. Moreover, these values represent the ‘full’ DM, as depicted by Eq. 2, including the actual MW and IGM foregrounds in addition to the host-galaxy contribution. We can subtract an estimated Milky Way contribution as provided in the TNS FRB report, i.e., . These FRBs have been localized to galaxies with measured redshifts too, and so we are able to subtract an estimated mean IGM contribution, using Eq. 3 and the fiducial cosmological parameters assumed in the Discussion. The resulting inferred observed host-galaxy FRB DMs are thus of the form: DM. Generally, we represent these observations in our Figures with a shaded band covering the 5-95% data range for the observed DMs, and show the median value (82.3, 255, & 156 pc cm-3, respectively). We do not break out the observed values by redshift, and compare them against all simulated redshifts as a means to discuss the general constraints of host-galaxy DMs at present.
@@ 3 Analysis
Observed FRBs | 82.3 | 156 | 201 | 255 | |
(DMObs-DMMW- |
Figure 1 shows maps of the free-electron column (i.e., dispersion measure, DM) on 100 pc scales in the six simulated galaxies analyzed here at , where we have summed the entire free-electron column 15 kpc444We did not sum the full kpc analysis volume purely for visualization reasons. Quantitatively, there is little to no difference as the column of ionized gas with is very low relative to the contribution from the innermost CGM () and full ISM ( kpc). above and below the disk mid-plane in each simulation (additional figures in Appendix A show DM maps of these simulated galaxies at and at the same spatial resolution). HII regions, easily visible as compact high-DM areas, are distributed across the galaxy disks and are indicative in their number of the general level of star formation in each simulation: m12r, being less massive, has generally lower levels of star formation and fewer HII regions; and m12w has a dramatically more compact gas disk, which is misaligned relative to the angular momentum of the stellar disk due to a recent LMC-mass merger, resulting in high overall DMs across its face. All of the disks broadly host full-column DMs on the order of 100-1000 pc cm-3 in their star-forming disks.
Focusing on the DMs measured to each star particle (ray-tracing from the particle to a height 20 kpc above the galaxy mid-plane), rather than the full free-electron column through the galaxy, we generate PDFs for the DM distribution towards all the star particles in all six galaxies at & 2 broken out for young (ages Myr) and old (ages Myr) populations, shown in Figure 2. These are the intrinsic DMs, i.e., not including any cosmological redshift term. In Appendix A, we present the DM distribution in each individual galaxy at & 2. We compare the DM distributions from the FIRE galaxies with the observed inferred FRB host-galaxy DMs (the total DMs less the MW-foreground and mean IGM estimates). Unsurprisingly, young stars (ages Myr) in the simulations produce hot, ionized HII regions around themselves which cause their DM distribution to be shifted to higher values. Broadly, there is significant overlap between the FIRE-2 host-galaxy DMs at all redshift and the observed FRB DM range (more so with the young-star DM distributions). Given that the observed DMs are presumed to be just DMHost, significant overlap of FIRE DMHost distributions allow us to infer possible FRB progenitor objects. We also include the outline of the orientation-averaged host-galaxy DM distribution for the snapshots, as those are the only ones that host disks (and thus are sensitive to inclination effects). Inclination on-average increases the host-galaxy DM distributions by up to a factor of 2, the effects are explored in more detail in § 3.1.
@@
Comparing to host-galaxy DM predictions from IllustrisTNG by Zhang et al. (2020), we see that the multiphase ISM in the FIRE-2 simulations hosts a far broader range of host-galaxy DMs than the much smoother effective equation of state ISM in the cosmological simulations. The offset in FIRE to higher median DMs relative to their ‘unbiased’ and ‘star-forming’ DM lines of sight, and the much broader scatter to high (and low) DMs in the wings of the distribution, near and demonstrates the importance of modeling a highly dynamic multiphase ISM in understanding the environments in which possible FRB progenitors live and evolve.
@@
Figure 3 shows both the median intrinsic (i.e., not incorporating for cosmological expansion) host-galaxy dispersion measure as a function of age for the star particles in the FIRE-2 galaxies at 0, 1, & 2 in ‘face-on’ orientations (the effects of inclination on DMs in the FIRE disks are explored in § 3.1, and we present the full stacked distribution of DM–stellar ages for each redshift in Appendix B). We directly compare the FIRE-2 host-galaxy dispersion measures with the observed inferred FRB population host-galaxy dispersion measure distribution (showing the median and 5-95% range). We add vertical lines in age at 3.4 and 37.4 Myr to denote the range of possible stellar lifetimes for core-collapse SN progenitors (specifically, the age range used in the FIRE-2 core-collapse feedback model). Across redshift, the FIRE-2 Milky Way progenitors show qualitatively similar distributions in dispersion measure as a function of age, despite growing considerably in mass and undergoing dramatic changes to their morphologies (i.e., settling into disks between and in this sample). Though stellar populations older than 10 Myr have median DMs roughly a 0.5-1 dex lower than the observed inferred FRB host-galaxy DMs, young populations (with ages Myr) have significantly higher median DMs, owing to their tendency to still be embedded in HII regions. Generally, the DMs begin falling before the onset of core-collapse SN feedback in FIRE, showing that pre-SN feedback (i.e., stellar jets, winds, and radiation) more than capable of dispersing the ionized gas regions surrounding young massive stars (see, e.g., Grudić et al., 2018; Lancaster et al., 2021; Grudić et al., 2022).
@@
Notably, m12w, a heavily disrupted disk galaxy having just underwent a major merger-driven starburst at with a highly compact gas disk, too hosts a stellar population with median host-galaxy dispersion measures sufficient to explain observed FRB DMs at nearly all stellar ages. This rules out building an assumption that only the lower-mass, star-bursting, high-gas fraction progenitors can produce such local ISM conditions.
@@ 3.1.1 Orientation-averaged Stacked DM-host PDFs
To further explore the effects of inclination, we stack together the different orientations of the FIRE disks at , weighting by their relative solid angle on (taking the inclinations , 30∘, 60∘, 90∘ and calculating the area of the bands between them). Figure 8 shows the result, for both the distribution of host-galaxy DMs as a function of stellar age (white line showing the weighted median), and the PDFs themselves, compared against inferred observed host-galaxy FRB DMs, and those calculated in the IllustrisTNG cosmological simulations by Zhang et al. (2020). We again find that the PDFs of host-galaxy DM are wider than previously appreciated, and that younger stars (10 Myr old) appear to agree best with the observationally inferred when the effects of orientation (inclination) are averaged over.
@@ 4 Discussion
Considering the expected mean IGM dispersion measure, as Yang et al. (2017) do to second order, we find:
where we have adopted the standard Planck Collaboration et al. (2016) cosmology ( km/s, , ), (see Yang et al. 2017, where we and they both assume full hydrogen and helium ionization in the IGM below ), (Shull et al., 2012), and , and are the speed of light, gravitational constant, and proton mass, respectively. Taking the ‘typical’ (median) DM excess (i.e., host-galaxy DM + IGM DM) from the observations of 391 pc cm-3, the ‘floor’ of 100 pc cm from the FIRE spirals implies a maximum redshift of (considering the remainder of the excess having come from the IGM), whereas if the typical FRB progenitor comes from a system more akin to one of the non-disky snapshots with a mean ‘young’ host-galaxy DM of 285 pc cm then the implied maximum redshift is only . For this calculation here we include the cosmological term for the host-galaxy contribution to the total DM. The general concern being that if the host-galaxy contribution to the observed DM excesses is larger than typically believed by factors of potentially a few, then the ability of FRBs to constrain cosmological models might be statistically significantly more limited than previously thought.
@@ 5 Conclusions
The median DMs we find in FIRE-2 galaxies are 2-60 times larger than the typical DMs through the effective equation of state ISMs of pc cm-3 found in the Illustris cosmological simulations and 1-7 times larger than those found in the IllustrisTNG cosmological simulations (Zhang et al., 2020), especially when considering inclination effects in disk galaxies.

### v2

@@ Abstract
We investigate the contribution of host galaxies to the overall Dispersion Measures (DMs) for Fast Radio Bursts (FRBs) using the Feedback in Realistic Environments (FIRE-2) cosmological zoom-in simulation suite. We calculate DMs from every star particle in the simulated L* galaxies by ray-tracing through their multi-phase interstellar medium (ISM), summing the line-of-sight free thermal electron column for all gas elements within 20 kpc of the galactic mid-plane. At , we find average (median) host-galaxy DMs of 74 (43) and 210 (94) pc cm-3 for older (10 Myr) and younger (10 Myr) stellar populations, respectively. Inclination raises the median DM measured for older populations (10 Myr) in the simulations by a factor of 2, but generally does not affect the younger stars deeply embedded in HII regions except in extreme edge-on cases (inclination ). In kinematically disturbed snapshots ( in FIRE), the average (median) host-galaxy DMs are higher: 80 (107) and 266 (795) pc cm-3 for older (10 Myr) and younger (10 Myr) stellar populations, respectively. FIRE galaxies tend to have higher DM values than cosmological simulations such as IllustrisTNG, with larger tails in their distributions to high DMs. As a result, FRB host galaxies may be closer (lower redshift) than previously inferred. Furthermore, constraining host-galaxy DM distributions may help significantly constrain FRB progenitor models.
@@ 1 Introduction
First discovered by Lorimer et al. (2007), Fast Radio Bursts (FRBs) are a mysterious class of millisecond duration radio transients that are likely extragalactic in origin. To date, hundreds of FRBs have been identified and a few tens have localized redshifts through the combined efforts of multiple teams using ASKAP, Parks, MOST, CHIME, DSA-110, and FAST telescopes (e.g., Thornton et al., 2013; Caleb et al., 2017; Bhandari et al., 2018; Shannon et al., 2018; CHIME/FRB Collaboration et al., 2019a; Heintz et al., 2020a; Niu et al., 2021; The CHIME/FRB Collaboration et al., 2021; Law et al., 2023). Additionally, some FRBs are known to be repeaters, providing some clues about their origins (e.g., Spitler et al., 2016; CHIME/FRB Collaboration et al., 2019a, b; Fonseca et al., 2020; Bhardwaj et al., 2021a). Despite this observational progress, the origins of FRBs remain mysterious (Cordes & Chatterjee, 2019; Petroff et al., 2019; Platts et al., 2019; Zhang, 2020; Petroff et al., 2021)111So much so that popular culture has picked up on the scientific debate, e.g., see: https://xkcd.com/2886/. One promising model for FRBs is that they are associated with magnetars; an idea which is strengthened by the detection of an FRB associated with the Milky Way (MW) magnetar SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020; Lu et al., 2020; Margalit et al., 2020). However, given its low energy compared to other FRBs, other source objects are certainly still viable (Cordes & Chatterjee, 2019; Petroff et al., 2019; Platts et al., 2019; Zhang, 2020). Ultimately, to determine the origins and nature of FRBs, their host galaxy and local environmental conditions must be understood, which requires more precise localization and understanding of the signal characteristics (Li et al., 2023; Mo et al., 2023).
@@
†Star formation rates averaged over a 10 Myr period, ignoring stellar evolutionary mass loss. |
@@
However, there is some evidence that FRBs have significantly larger host-galaxy DMs than suggested by cosmological simulations. For example, repeating FRB 20190520B has an estimated host-galaxy pc cm-3 (after accounting for foreground galaxy clusters) and a redshift of (Niu et al., 2022a; Ocker et al., 2022; Anna-Thomas et al., 2022; Lee et al., 2023a), which provides a cautionary example of interpreting high-DM FRBs as having high redshifts.
@@ 2.2 Observed FRB Dispersion Measures
To compare our calculated host-galaxy DMs against the observed distribution of FRB DMs, we searched for reported FRB events: as of July 31, 2024, 43 FRBs with measured DMs (), localized redshifts with host galaxy identifications, and estimates for Milky Way foreground DM () were found in the literature, which we took as a representative population. Appendix A lists the details of this FRB sample, including details of the host galaxies and localization references. As well, the redshift and dispersion measure (minus MW contributions) distributions can be found therein.
For consistency in our comparison, we have taken the reported ‘full’ DM values (i.e., that of Eq. 2) as reported in the literature for each of the FRBs. We subtract the estimated Milky Way contribution as provided by the source (generally this has been from the NE2001 model of Cordes & Lazio 2002), i.e., . These FRBs have been localized to galaxies with measured redshifts too, and so we are able to subtract an estimated mean IGM contribution following Yang et al. (2017):
where we have adopted the standard Planck Collaboration et al. (2016) cosmology ( km/s, , ), (see Yang et al. 2017, where we and they both assume full hydrogen and helium ionization in the IGM below ), (Shull et al., 2012), and , and are the speed of light, gravitational constant, and proton mass, respectively. The resulting inferred observed host-galaxy FRB DMs are thus of the form: DM. Generally, we represent these observations in our Figures with a shaded band covering the 5-95% data range for the observed DMs, and show the median value (55.1, 545, & 157.6 pc cm-3, respectively). Except for one observed FRB, we do not break out the observed values by redshift, and compare them against all simulated redshifts as a means to discuss the general constraints of host-galaxy DMs at present. The only exception being FRB 20220610A reported by Gordon et al. (2024), having been localized to a star-forming galaxy at , which we compare directly against the FIRE-2 simulations at . For the purposes of comparing this FRB, we plot their predicted pc cm-3 directly (not recalculating their reported mean IGM contribution, and the resulting shift in DMHost).
@@ 3 Analysis
Host galaxy-localized FRBs | 55.1 | 157.6 | 193.8 | 545.0 | |
(DMObs-DMMW-) |
Figure 1 shows maps of the free-electron column (i.e., dispersion measure, DM) on 100 pc scales in the six simulated galaxies analyzed here at , where we have summed the entire free-electron column 15 kpc444We did not sum the full kpc analysis volume purely for visualization reasons. Quantitatively, there is little to no difference as the column of ionized gas with is very low relative to the contribution from the innermost CGM () and full ISM ( kpc). above and below the disk mid-plane in each simulation (additional figures in Appendix B show DM maps of these simulated galaxies at and at the same spatial resolution). HII regions, easily visible as compact high-DM areas, are distributed across the galaxy disks and are indicative in their number of the general level of star formation in each simulation: m12r, being less massive, has generally lower levels of star formation and fewer HII regions; and m12w has a dramatically more compact gas disk, which is misaligned relative to the angular momentum of the stellar disk due to a recent LMC-mass merger, resulting in high overall DMs across its face. All of the disks broadly host full-column DMs on the order of 100-1000 pc cm-3 in their star-forming disks.
Focusing on the DMs measured to each star particle (ray-tracing from the particle to a height 20 kpc above the galaxy mid-plane), rather than the full free-electron column through the galaxy, we generate PDFs for the DM distribution towards all the star particles in all six galaxies at & 2 broken out for young (ages Myr) and old (ages Myr) populations, shown in Figure 2. In Appendix B, we present the DM distribution in each individual galaxy at & 2. We compare the DM distributions from the FIRE galaxies with the observed inferred FRB host-galaxy DMs (the total DMs less the MW-foreground and mean IGM estimates). Unsurprisingly, young stars (ages Myr) in the simulations produce hot, ionized HII regions around themselves which cause their DM distribution to be shifted to higher values. Broadly, there is significant overlap between the FIRE-2 host-galaxy DMs at all redshift and the observed FRB DM range (more so with the young-star DM distributions). Given that the observed DMs are presumed to be just DMHost, significant overlap of FIRE DMHost distributions allow us to infer possible FRB progenitor objects. We also include the outline of the orientation-averaged host-galaxy DM distribution for the snapshots, as those are the only ones that host disks (and thus are sensitive to inclination effects). Inclination on-average increases the host-galaxy DM distributions by up to a factor of 2, the effects are explored in more detail in § 3.1.
@@
Comparing to host-galaxy DM predictions from IllustrisTNG by Mo et al. (2023), we see that the multiphase ISM in the FIRE-2 simulations hosts a far broader range of host-galaxy DMs than the much smoother effective equation of state ISM in the cosmological simulations. We have subtracted their reported CGM contributions to DMhost, of 17% and 40% for their TNG-SFR and TNG-M⋆, respectively. The offset in FIRE to higher median DMs relative to their star-forming host galaxy DM (“TNG-SFR”, with specific SFRs ( yr-1) and stellar mass-weighted host galaxy DM (“TNG-M⋆”, weighted by stellar mass in their sample) lines of sight, and the much broader scatter to high (and low) DMs in the wings of the distribution, near and demonstrates the importance of modeling a highly dynamic multiphase ISM in understanding the environments in which possible FRB progenitors live and evolve.
@@
Figure 3 shows both the median host-galaxy dispersion measure as a function of age for the star particles in the FIRE-2 galaxies at 0, 1, & 2 in ‘face-on’ orientations (the effects of inclination on DMs in the FIRE disks are explored in § 3.1, and we present the full stacked distribution of DM–stellar ages for each redshift in Appendix C). We directly compare the FIRE-2 host-galaxy dispersion measures with the observed inferred FRB population host-galaxy dispersion measure distribution (showing the median and 5-95% range). We add vertical lines in age at 3.4 and 37.4 Myr to denote the range of possible stellar lifetimes for core-collapse SN progenitors (specifically, the age range used in the FIRE-2 core-collapse feedback model). Across redshift, the FIRE-2 Milky Way progenitors show qualitatively similar distributions in dispersion measure as a function of age, despite growing considerably in mass and undergoing dramatic changes to their morphologies (i.e., settling into disks between and in this sample). Though stellar populations older than 10 Myr have median DMs roughly a 0.5-1 dex lower than the observed inferred FRB host-galaxy DMs, young populations (with ages Myr) have significantly higher median DMs, owing to their tendency to still be embedded in HII regions. Generally, the DMs begin falling before the onset of core-collapse SN feedback in FIRE, showing that pre-SN feedback (i.e., stellar jets, winds, and radiation) more than capable of dispersing the ionized gas regions surrounding young massive stars (see, e.g., Grudić et al., 2018; Lancaster et al., 2021; Grudić et al., 2022).
@@
In the and snapshots, given their irregular morphology and dispersion supported nature, we clearly see in the DM maps (Figs. 10 & 11) that they are dominated by a few large star-forming regions (rather than the distributed star formation seen in the disks). Consequently, the host DMs measured for the youngest stellar populations correspond directly, or nearly so, to the recent SFR of the entire galaxy (cf. Table 1), as the global intensity of the star formation is much more tightly related to the radiation field seen by the gas around the youngest stars. Whereas, the distributed nature of star formation in the disks at means that the global star formation rate is not as important in setting the local radiation field seen by gas around young stars, and the individual star forming regions are significantly smaller producing smaller ionized gas columns around young stars/star-forming regions.
Notably, m12w, a heavily disrupted disk galaxy having just underwent a major merger-driven starburst at with a highly compact gas disk, too hosts a stellar population with median host-galaxy dispersion measures sufficient to explain observed FRB DMs at nearly all stellar ages. This challenges the assumption that only the lower-mass, star-bursting, high-gas fraction progenitors can produce such local ISM conditions.
@@ 3.1.1 Orientation-averaged Stacked DM-host PDFs
To further explore the effects of inclination, we stack together the different orientations of the FIRE disks at , weighting by their relative solid angle on (taking the inclinations , 30∘, 60∘, 90∘ and calculating the area of the bands between them). Figure 8 shows the result, for both the distribution of host-galaxy DMs as a function of stellar age (white line showing the weighted median), and the PDFs themselves, compared against inferred observed host-galaxy FRB DMs, and those calculated in the IllustrisTNG cosmological simulations by Mo et al. (2023). We again find that the PDFs of host-galaxy DM are wider than previously appreciated, and that younger stars (10 Myr old) appear to agree best with the observationally inferred when the effects of orientation (inclination) are averaged over.
@@ 4 Discussion
Considering the expected mean IGM dispersion measure, as Yang et al. (2017) also do to second order (Eq. 3 being the exact integral), we find:
where we have adopted the standard Planck Collaboration et al. (2016) cosmology ( km/s, , ), (see Yang et al. 2017, where we and they both assume full hydrogen and helium ionization in the IGM below ), and (Shull et al., 2012). Taking the ‘typical’ (median) DM excess (i.e., host-galaxy DM + IGM DM) from the observations of 391 pc cm-3, the ‘floor’ of 100 pc cm from the FIRE spirals implies a maximum redshift of (considering the remainder of the excess having come from the IGM), whereas if the typical FRB progenitor comes from a system more akin to one of the non-disky snapshots with a mean ‘young’ host-galaxy DM of 285 pc cm then the implied maximum redshift is only . For this calculation here we include the cosmological term for the host-galaxy contribution to the total DM. The general concern being that if the host-galaxy contribution to the observed DM excesses is larger than typically believed by factors of potentially a few, then the ability of FRBs to constrain cosmological models might be statistically significantly more limited than previously thought.
@@ 5 Conclusions
The median DMs we find in FIRE-2 galaxies are 2-60 tim larger than the typical DMs through the effective equation of state ISMs of pc cm-3 found in the Illustris cosmological simulations and 1-2 times larger than those found in the IllustrisTNG cosmological simulations (Zhang et al., 2020; Mo et al., 2023), especially when considering inclination effects in disk galaxies.
@@
Appendix A Compiled host galaxy-localized FRB Observations
FRB ID | DM - DM | Redshift | Host Galaxy | Notes | SFR | Localization | |
[pc cm-3] | [M⊙] | [M⊙/yr] | Reference(s) | ||||
20221012A | 386.68 | 0.284669 | PSO J280.8014+70.5242 | Quiescent | 10.99 | 0.15 | (1) |
20220920A | 274.69 | 0.158239 | PSO J240.2568+70.9180 | 9.81 | 4.10 | (1) | |
20220914A | 576.08 | 0.113900 | PSO J282.0585+73.3363 | 9.48 | 0.72 | (1) | |
20220825A | 571.54 | 0.241397 | PSO J311.9820+72.5848 | 9.95 | 1.62 | (1) | |
20220610A | 1320† | 1.017 | High- star-forming galaxy | 9.7 | 1.7 | (2) | |
20220509G | 214.33 | 0.089400 | PSO J282.6748+70.2427 | Spiral | 10.79 | 0.23 | (1) |
20220506D | 307.87 | 0.30039 | PSO J318.0447+72.8272 | 10.29 | 5.08 | (1) | |
20220418A | 585.65 | 0.622000 | PSO J219.1065+70.0953 | 10.31 | 29.41 | (1) | |
20220310F | 416.84 | 0.477958 | PSO J134.7211+73.4910 | 9.97 | 4.25 | (1) | |
20220307B | 363.57 | 0.248123 | PSO J350.8747+72.1918 | 10.04 | 2.71 | (1) | |
20220207C | 183.08 | 0.043040 | PSO J310.1977+72.8826 | Star-forming disk galaxy | 9.91 | 1.16 | (1) |
20211212A | 178.9 | 0.0715 | Star-forming spiral | 10.28 | 0.73 | (3), (4), (5) | |
20211203C | 572.8 | 0.3439 | Star-forming | 9.76 | 15.91 | (3), (4), (6) | |
20211127I | 192.33 | 0.0469 | WALLABY J131913-185018 | Star-forming spiral | 9.48 | 35.83 | (7), (5) |
20210807D | 130.7 | 0.1293 | Quiescent spiral | 10.97 | 0.63 | (3), (5) | |
20210410D | 522.58 | 0.1415 | J214420.69-791904.8 | Star-forming | 9.47 | 0.03 | (8) |
20210405I | 50.3 | 0.066 | J1701249-4932475 | Star-forming spiral | 11.25 | 0.3 | (9) |
20210320C | 342.6 | 0.2797 | Star-forming | 10.37 | 3.51 | (3), (4), (6) | |
20210117A | 695.1 | 0.2145 | J223955.07-160905.37 | Dwarf galaxy | 8.59 | 0.02 | (10) |
20201124A | 287.03 | 0.0979 | SDSS J050803.48+260338.0 | Massive dusty star-forming galaxy | 10.23 | 2.1 | (11), (12), (13) |
20200906A | 541.8 | 0.3688 | PSO J033358.994-140459.287 | Star-forming galaxy | 10.12 | 0.48 | (14) |
20200430A | 352.9 | 0.1608 | PSO J151849.52+122235.8 | No reported structure | 9.11 | 0.2 | (15) |
20200223B | 156.2 | 0.06024 | SDSS J003304.68+284952.6 | 10.75 | 0.59 | (16) | |
20191228A | 264.5 | 0.2432 | Star-forming galaxy | 9.73 | 0.50 | (14) | |
20191106C | 307.2 | 0.10775 | SDSS J131819.23+425958.9 | 10.65 | 4.75 | (16) | |
20191001A | 463.9 | 0.2340 | DES J213324.44-544454.65 | Star-forming spiral, merger | 10.7 | 11.2 | (17), (18), (19) |
20190714 | 466.13 | 0.2365 | PSO J121555.0941-130116.004 | Possible spiral arms | 10.17 | 0.65 | (15) |
20190711A | 537.23 | 0.5217 | No structure | 8.91 | 0.42 | (15) | |
20190611 | 263.57 | 0.3778 | J212258.0-792350 | No spiral arms | 8.9 | 0.27 | (15) |
20190608B | 301.5 | 0.1178 | SDSS J221604.90-075355.9 | Star-forming spiral | 10.4 | 1.2 | (20) |
20190523A | 723.8 | 0.66 | DSA-10 J134815.6+722811 | Quiescent | 10.79 | 0.09 | (15) |
20190520B | 642∗ | 0.241 | J160204.31-111718.5 | Dwarf galaxy | 8.78 | 0.41 | (21), (22) |
20190425A | 79.2 | 0.03122 | UGC 10667 | Star-forming spiral | 10.26 | 1.5 | (23) |
20190418A | 113.5 | 0.07132 | SDSS J042314.96+160425.6 | Star-forming spiral | 10.27 | 0.15 | (23) |
20190102C | 306.3 | 0.291 | ASKAP-ICS J212939.76-792832.5 | Star-forming | 9.53 | 0.86 | (15) |
20181223C | 92.5 | 0.03024 | SDSS J120340.98+273251.4 | Star-forming spiral | 9.29 | 0.054 | (23) |
20181220A | 83.4 | 0.02746 | 2MFGC 17440 | Star-forming spiral | 9.86 | 3.0 | (23) |
20181112A | 487.27 | 0.4755 | ASKAP-ICS J214923.63-525815.4 | Star-forming, no identified spiral | 9.42 | 0.6 | (20) |
20181030A | 62.5 | 0.0039 | NGC 3252 | Star-forming barred spiral galaxy | 9.76 | 0.355 | (24) |
20180924B | 320.92 | 0.3214 | ASKAP-ICS J214425.255-405400.1 | Star-forming | 10.35 | 0.88 | (15), (20) |
20180916B | 149.06 | 0.0337 | SDSS J015800.28+654253.0 | Star-forming/Quiescent spiral | 9.33 | 0.06 | (15) |
20180301A | 384 | 0.3304 | PSO J093.2268+04.6703 | Star-forming galaxy | 9.36 | 1.93 | (14) |
20121102A | 372.58 | 0.19273 | J053158.698+330852.60 | Star-forming dwarf galaxy | 8.14 | 0.15 | (25) |
Note. — DMs all in units of pc cm-3, all values in parentheses include 1/(1+) cosmological redshift term.
References. — (1) Law et al. 2024, (2) Gordon et al. 2024, (3) Gordon et al. 2023, (4) James et al. 2022, (5) Deller in prep., (6) Shannon in prep., (7) Glowacki et al. 2023, (8) Caleb et al. 2023, (9) Driessen et al. 2024, (10) Bhandari et al. 2023, (11) Ravi et al. 2022, (12) Fong et al. 2021, (13) Piro et al. 2021, (14) Bhandari et al. 2022, (15) Heintz et al. 2020b, (16) Ibik et al. 2024, (17) Bhandari et al. 2020a, (18) Kundu 2022, (19) Woodland et al. 2023, (20) Bhandari et al. 2020b, (21) Niu et al. 2022b, (22) Lee et al. 2023b, (23) Bhardwaj et al. 2023, (24) Bhardwaj et al. 2021b, (25) Tendulkar et al. 2017.
We have assembled, as far as the authors can tell, a near complete collection of the published FRBs with identified host galaxies and redshifts as of July 31, 2024 to compare against the host galaxy dispersion measure distributions found in the FIRE-2 galaxies. The FRBs, their identifiers, extragalactic DMs, redshifts, host galaxy identifiers, stellar masses, star formation rates, and localization references can be found in Table A. Two of the FRBs, 20190520B and 20220610A have additional reported DM contributions from galaxy clusters/groups along the line of sight, which we have subtracted according to their respective source publications. A large fraction of the reported host galaxies, where morphology has been noted do appear to be star-forming spirals (about a third of the sample, though by-and-large the whole sample of host galaxies has not been observed at high spatial resolution). Figure 9 shows both the inferred host galaxy DM distribution (by subtracting the mean IGM component, as described in 2), from which we derive our 5-95% observations distribution that we compare against FIRE-2 in figures in the main text, and the extragalactic DM excess–redshift (the ‘Macquart et al. 2020’ relation) showing the derived 5-95% range of inferred .